### v1

# The role of equation of state on the spin-up of millisecond pulsars

Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength. The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS. Finding more MSPs with low mass and and short spin-period could help to put more strict constraints on the EoS of pulsars.
dense matter — stars: neutron — pulsars: genera
1 Introduction
The pulsar-like compact stars are idea laboratories to explore the state of matter at supra-nuclear densities. The equation of state (EoS) of pulsars is still a controversial topic at present, although abundant observational data of pulsars has been accumulated. Based on different points of view, a variety models for pulsars have been speculated. The neutron star (NS) model whose basic idea could be found in Landau (1932) and the quark star (QS) model which originated from Witten’s conjecture (Bodmer, 1971; Witten, 1984) have attracted most attention. The strangeon star (SS) model was also proposed (Xu, 2003) to understand different manifestations of pulsars, where the strong coupling between quarks has grouped quarks into strangeons (i.e. “strange nucleons”) , and has passed various observational tests (see Lai et al. (2023) and references therein).
The EoS of pulsars is essentially depends on the behavior of quantum chromodynamics (QCD) at low-energy scales, which is still a challenge for us to understand. Although the constraints by both the existence of high-mass pulsars and the tidal deformability of GW 170817 (Abbott et al., 2017) could dramatically reduce the allowed EoSs of NSs (Annala et al., 2018), it is still uncertain that whether pulsars are NSs, QSs or SSs. In addition, these three kinds of models mentioned above have subclass models respectively, e.g., some branches of NS model still allow the existence of quark matter at the core region (Weber, 2005), QS model includes different states such us unpaired strange quark matter state (Alcock et al., 1986) and colour super-conductivity (CSC) state (Alford et al., 2008), and the hybrid SS model is proposed (Zhang et al., 2023) where an SS develops a strange quark matter core. Therefore, the differences among models could be subtle. It is certainly important to find out unambiguous and observable probes to tell what is the real nature of pulsars.
The global structure of pulsars, e.g. the mass and radius, can provide robust constraints on EoS models when compared with observations (Lattimer & Prakash, 2007; Özel et al., 2010). Despite the uncertainty about the explicit structure determined by various classes of EoS models, there are some general differences in the global structures between NSs and QSs/SSs (Demorest et al., 2010; Gao et al., 2022). NSs are gravity-bound so that the surface density could approach zero, and the radii of low-mass NSs (below $\sim M_{\odot}$) decrease with increasing masses and the radii of intermediate-mass NSs (in the range between $\sim M_{\odot}$ and $\sim 2M_{\odot}$) could remain nearly constant. Both QSs and SSs are self-bound by strong interaction so that the surface density could still above the nuclear matter density, and their radii increase with increasing masses unless the mass approaches its maximum value. The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.5M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not. Therefore, testing whether pulsars are gravity-bound or self-bound provides a relatively less unambiguous probe to the nature of pulsars.
In this paper we investigate that how different global structures of pulsars from different EoS models affect the spin-up process of millisecond pulsars (MSPs). MSPs evolve from pulsars in binaries which have underwent the spin-up process, i.e. recycling precess, due to accreting mass from their companions. The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000, 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star. The observed differences of the spin distributions of recycled pulsars with different types of companions have been explained by the amount of accreted mass needed for a pulsar to achieve its equilibrium spin (Tauris et al., 2012). In addition, the magnetic field evolution of accreting pulsars would also affect the formation of MSPs. The spin-period evolutions of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006). The minimum spin-period of millisecond pulsars also depends the correlation between the accretion rate and the surface dipole magnetic field in LMXBs (Ertan & Alpar, 2021).
The spin-up line in the $P\dot{P}$-diagram is often used to demonstrate the formation and evolution of recycled pulsars, although it in fact cannot be uniquely defined (Tauris et al., 2012) due to the dependence on both the process of accretion and the properties of the pulsar. Most of the previous studies about the evolution of MSPs focusing on the accretion process in the binary evolution use the typical values for the structure of neutron stars. These studies discuss many parameters which do affect the location of the spin-up line in the $P\dot{P}$-diagram, e.g. the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979a). It is worth noting that, the spin evolution of a pulsar also depends on the EoS which determines the pulsar’s global structure such as mass, radius and the moment of inertia. In this paper we will investigate the role of equation of state on the spin-up of MSPs.
Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs. Combining the spin-up condition derived in Sec.2 and the minimal spin-period of a recycled pulsar estimated in Sec.3, we can constrain the upper limit for the magnetic field and the lower limit for the accreted mass. The results shown in Sec.4 indicate that it could be easier for an SS to form a low-mass ($M<1.5M_{\odot}$) and fully recycled ($P<10$ ms) MSP than that for an NS. Conclusions and discussions are made in Sec.5.
2 The spin-up line
[…]
Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979a), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter. To exert a positive accretion torque, $\omega_{\rm c}$ is about 0.2-1 (Ghosh & Lamb, 1979b; Wang et al., 2011).
The spin-up condition can therefore be expressed as the relation between $r_{\rm A}$ and $r_{\rm co}$, where the corotation radius $r_{\rm co}=[GMP^{2}/(4\pi^{2})]^{1/3}$ is defined as the radius where the Keplerian frequency matches the spin frequency of the pulsar. Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as $r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}$. Because we will focus on the EoS-dependent results of the spin-up process, the dependence on other factors, such as truncation of the accretion disc and the magnetic field configuration outside the star, will not be included when we are comparing different EoS models. Taking into account the possible values for the factors $\phi$ and $\omega_{\rm c}$ mentioned before, we assume $\omega_{\rm c}^{2/3}/\phi=0.5$, so the spin-up condition is
$r_{\rm A}\leq 0.5\ r_{\rm co}.$ (1)
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities. Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram, which would give the maximum allowed magnetic field $B_{\rm max}$ for the further spin-up process.
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram. and Fig.1 shows the results for NS and SS models. Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively. We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass. The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
Figure 1: The spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
3 The minimal spin-period by accretion
By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by ${\rm d}J=n\xi\sqrt{GMr_{\rm A}}{\rm d}M$, where $\xi\simeq 1$ is a numerical factor depending on the flow pattern (Ghosh & Lamb, 1979a) and $n$ is a dimensionless torque which could be assumed to be $\simeq 1$ (Tauris et al., 2012). Then the final spin-period $P$ of a recycled pulsar (from rest) from the initial mass $M_{1}$ to the final mass $M_{2}$ could be estimated by
$P^{-1}_{\rm min}=\frac{1}{2\pi I}\int_{M_{1}}^{M_{2}}\sqrt{GMr_{\rm A}}{\rm d}M,$ (2)
where $I$ is the (final) moment of inertia. We have assumed above that all the orbital angular momentum of the accreted material is transferred to the spin angular momentum of the pulsar, so $P$ in equation (2) is the minimal value of the final spin-period. In the following the values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms.
Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables. To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumptions as in Tauris et al. (2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium. In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak. Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$. Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
The $P_{\rm min}$ and $M$ (final mass) relation for SS and NS models are shown in Fig.2, denoted respectively by red lines and blue lines, where solid lines represent the case of $\Delta M=0.05M_{\odot}$ and dashed lines represent the case of $\Delta M=0.1M_{\odot}$. We can see that, with the same amount of accreted mass, NSs would spin faster than SSs. This is understandable: the radius of an NS is usually larger than that of an SS with the same mass, so an NS usually has larger $r_{\rm A}$ than an SSs, which means that an NS could gain more angular momentum than an SS when the accreted mass is the same. In Fig.2 we choose $B=5\times 10^{8}$ G, and smaller values of $B$ will lift the lines upwards, since smaller $r_{\rm A}$ will lead to smaller angular momentum transfer.
Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G. The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively. The horizontal ordinate denotes the final mass after accretion.
4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$. Because equation (2) gives the minimal value of the final spin-period $P_{\rm min}$ for a given value of $\Delta M$, for a pulsar with a known value of $P$ it gives the minimum accreted mass $\Delta M_{\rm min}$. In the $MP$ diagram, where $M$ denotes the final mass after accretion, we can plot the contours of $B_{\rm max}$ for a given EoS by combining equations (1) and (2).
In the $MP$ diagram, a higher contour of $B_{\rm max}$ corresponds to a higher value of $B_{\rm max}$, and a lower curve of $\Delta M_{\rm min}$ corresponds to a higher value of $\Delta M_{\rm min}$. Therefore, if we could determine the lower limit for $B$ and upper limit for $\Delta M$, i.e. $B\geq B_{0}$ and $\Delta M\leq\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$. Fig.3 shows the allowed regions of SSs (left) and NSs (right), where we adopt $\Delta M_{\rm min}=0.05M_{\odot}$ for solid curves and $B_{\rm max}=5\times 10^{8}$ G for dashed curves. From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
Refer to captionRefer to caption
Figure 3: The allowed regions of SSs (left) and NSs (right), where solid curves denote $\Delta M_{\rm min}=0.05M_{\odot}$ and dashed curves denote $B_{\rm max}=5\times 10^{8}$ G. The horizontal ordinate denotes the final mass after accretion. If we could determine the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$.
As mentioned above, to set the allowed region in $MP$ diagram we should know the the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, although there are a lot of uncertainties to constrain $B$ and $\Delta M$. We could estimate $B_{0}$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, so by setting $\sin\alpha=1$ we can get $B_{0}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$. It is interesting to noted that, for each pulsar whose data of $M$ and $P$ is plotted in Fig.4 (see below), the value of $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue  [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/  (Manchester et al., 2005) is larger than the value of $B$ of the nearest dashed curve below it in $MP$ diagram, under both SS and NS models. So setting $B_{0}$ as the lower limit of $B$ seems to be a reasonable way. It is also possible to estimate $\Delta M_{0}$ from the evolution process of the binary system, such as the accretion rate, the age of the MSP and the interaction between the MSP and the donor star.
To compare with observations, we extend the allowed regions shown in Fig.3 to different values of $B_{\rm max}$ and $\Delta M_{\rm min}$, and we show in Fig.4 the results for SSs (left) and NSs (right). Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
Refer to captionRefer to caption
Figure 4: Contours of $\Delta M_{\rm min}$ and $B_{\rm max}$ for SSs (left) and NSs (right). The horizontal ordinate denotes the final mass after accretion. Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.4. Although the curves of $\Delta M_{\rm min}$ for SSs are not significantly different from that for NSs, the differences between the contours of $B_{\rm max}$ for SSs and that for NSs are visible. The dashed curves in Fig. 4 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs. The constraints on EoS models are more significant for masses below about $1.6M_{\odot}$ than that for higher masses. For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could. And on the other hand, if the magnetic field is higher than $8\times 10^{8}$ G, then an SS with mass approaching $2.2M_{\odot}$ could not be spun up to $P<4$ ms, but an NS could.
As more and more masses of MSPs with increasing accuracy have been being measured, we can study the recycling of pulsars by comparison with observations in $MP$ diagram. Although from the data plotted in Fig. 4 we could not yet disfavor neither SS nor NS models, finding more MSPs with $P<10$ ms with neither low mass or high mass in the future could help to put more strict constraints on the EoS of pulsars.
5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs. The allowed regions in $MP$ diagram derived by the lower limit for $B$ and the upper limit for $\Delta M$ indicate that, it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled MSPs with neither low mass (below about $1.5M_{\odot}$) or very high mass (above about $2.2M_{\odot}$) in the future could be helpful.
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs. It is worth adding the role of EoS to other effects, such as the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion, to study the recycling of MSPs more comprehensively.
The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions. A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Finding more MSPs with spin periods less than 10 milliseconds and with accurate mass-measurement in the future, e.g. by SKA and FAST, could help to put more strict constraints on the nature of pulsars.
This work is supported by the National SKA Program of China (No. 2020SKA0120300) and the Young Top-notch Talent Cultivation Program of Hubei Province.

### v2

@@
 Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength.
+The magnetic field strength will be important in impeding the spin-up of NSs, especially to the ones with low masses, but the impeding effect would not be important to SSs.
 The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS.
-Finding more MSPs with low mass and and short spin-period could help to put more strict constraints on the EoS of pulsars.
-dense matter — stars: neutron — pulsars: genera
+Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on magnetic field strength, could help to put more strict constraints on the EoS of pulsars.
+dense matter — stars: neutron — pulsars: general
 1 Introduction
@@
 Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs.
-Combining the spin-up condition derived in Sec.2 and the minimal spin-period of a recycled pulsar estimated in Sec.3, we can constrain the upper limit for the magnetic field and the lower limit for the accreted mass.
-The results shown in Sec.4 indicate that it could be easier for an SS to form a low-mass ($M<1.5M_{\odot}$) and fully recycled ($P<10$ ms) MSP than that for an NS.
+From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process.
+To take into account the effect of accretion process, in Sec.3 we derive the minimal spin-period by assuming that all the angular momentum of the accreted material is added the the pulsar.
+Although the results of Sec.3 show that NSs would spin faster than SSs with the same amount of accreted mass and the same magnetic field, the minimal-spin contours derived in Sec.4 show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same.
+The results shown in Sec.4 indicate that the constraints on EoS models are significant for masses below about $1.5M_{\odot}$, especially when we get more accurate estimation of the magnetic field strength.
 Conclusions and discussions are made in Sec.5.
@@
 To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities.
-Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram, which would give the maximum allowed magnetic field $B_{\rm max}$ for the further spin-up process.
+Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram.
 The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent.
-Different EoS models will give different spin-up lines in $MP$ diagram.
-and Fig.1 shows the results for NS and SS models.
+Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.1 shows the results for NS and SS models.
 Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively.
@@
 The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
-Figure 1: The spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
+Figure 1: Spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
 3 The minimal spin-period by accretion
@@
 Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables.
-To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumptions as in Tauris et al.
-(2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium.
+To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumption as in Tauris et al.
+(2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium (but this assumption can be relaxed in deriving the minimal-spin contour which contains different values of $B$, see Sec.4).
 In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak.
@@
 The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively.
-The horizontal ordinate denotes the final mass after accretion.
+In $MP$ diagram, the horizontal ordinate denotes the final mass after accretion.
 4 Comparison with observations
@@
 Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
-Because equation (2) gives the minimal value of the final spin-period $P_{\rm min}$ for a given value of $\Delta M$, for a pulsar with a known value of $P$ it gives the minimum accreted mass $\Delta M_{\rm min}$.
-In the $MP$ diagram, where $M$ denotes the final mass after accretion, we can plot the contours of $B_{\rm max}$ for a given EoS by combining equations (1) and (2).
-In the $MP$ diagram, a higher contour of $B_{\rm max}$ corresponds to a higher value of $B_{\rm max}$, and a lower curve of $\Delta M_{\rm min}$ corresponds to a higher value of $\Delta M_{\rm min}$.
-Therefore, if we could determine the lower limit for $B$ and upper limit for $\Delta M$, i.e.
-$B\geq B_{0}$ and $\Delta M\leq\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$.
-Fig.3 shows the allowed regions of SSs (left) and NSs (right), where we adopt $\Delta M_{\rm min}=0.05M_{\odot}$ for solid curves and $B_{\rm max}=5\times 10^{8}$ G for dashed curves.
+If the values of both $B$ and $\Delta M$ of a recycled pulsar is known, we can plot the allowed region in $MP$ diagram which satisfies both equations (1) and (2), i.e.
+by putting Fig.1 and Fig.2 together.
+We show the allowed regions of SSs (left) and NSs (right) for $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G in Fig.3, where dashed line denotes the general spin-up line (cf.
+Fig.1) and solid line shows the minimal spin-period by accretion (cf.
+Fig.2).
 From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
 Refer to captionRefer to caption
-Figure 3: The allowed regions of SSs (left) and NSs (right), where solid curves denote $\Delta M_{\rm min}=0.05M_{\odot}$ and dashed curves denote $B_{\rm max}=5\times 10^{8}$ G.
-The horizontal ordinate denotes the final mass after accretion.
-If we could determine the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$.
-As mentioned above, to set the allowed region in $MP$ diagram we should know the the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, although there are a lot of uncertainties to constrain $B$ and $\Delta M$.
-We could estimate $B_{0}$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, so by setting $\sin\alpha=1$ we can get $B_{0}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$.
-It is interesting to noted that, for each pulsar whose data of $M$ and $P$ is plotted in Fig.4 (see below), the value of $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue  [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/  (Manchester et al., 2005) is larger than the value of $B$ of the nearest dashed curve below it in $MP$ diagram, under both SS and NS models.
-So setting $B_{0}$ as the lower limit of $B$ seems to be a reasonable way.
-It is also possible to estimate $\Delta M_{0}$ from the evolution process of the binary system, such as the accretion rate, the age of the MSP and the interaction between the MSP and the donor star.
-To compare with observations, we extend the allowed regions shown in Fig.3 to different values of $B_{\rm max}$ and $\Delta M_{\rm min}$, and we show in Fig.4 the results for SSs (left) and NSs (right).
-Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
+Figure 3: The allowed regions of SSs (left) and NSs (right) in $MP$ diagram, where solid curves denote minimal-spin lines with $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G, and dashed curves denote spin-up lines with $B=5\times 10^{8}$ G.
+However, in most cases we are not certain about both the value of $B$ and $\Delta M$ for recycled pulsars.
+If we estimate $B$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, then by setting $\sin\alpha=1$ we can get $B>1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$.
+It is interesting to noted that, each pulsar whose data of $M$ and $P$ is plotted in Fig.5 (see below) is above the spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue (https://www.atnf.csiro.au/research/pulsar/psrcat/(Manchester et al., 2005) under both SS and NS models.
+As mentioned before, in this paper we keep $B$ as a parameter, instead of replacing in some model-dependent ways.
+Then how to show the difference between SSs and NSs models by comparing with observations?
+Combining equations (1) and (2) we can get the minimal-spin contour by accretion of $\Delta M$, which is defined as the contour of $\Delta M$ in $MP$ diagram, representing the minimal spin-period by accretion of $\Delta M$ for different values of $B$.
+For a data point in $MP$ diagram, there exists both a minimal-spin contour and a spin-up line intersecting it.
+The value of $\Delta M$ of the minimal-spin contour is the minimal accreted mass of the pulsar, while the value of $B$ of the spin-up line is the upper limit on its magnetic field strength.
+If we can use some other methods to constrain $\Delta M$ (e.g.
+from the binary evolution) and $B$ (e.g.
+theoretically from the magnetic field decay in accretion, or observationally from the cyclotron absorption lines), it would be hopeful to see that whether this pulsar is more likely to be an NS or an SS.
+From Fig.3 we can infer that low mass and fast spinning pulsars would present significant different parameters between SS and NS models, so we use a hypothetical data point in $MP$ diagram as an example in Fig.4.
+If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G.
+One the other hand, if this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
+This example shows that the difference in the the upper limit on $B$ between SS and NS models is significant for low mass and fast spinning pulsars.
+Figure 4: We use a hypothetical data point in $MP$ diagram as an example.
+If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G.
+If this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
+To compare with actual data, we extend the curves in Fig.4 to different values of $B$ and $\Delta M$, and we show in Fig.5 the results for SSs (left) and NSs (right).
+Solid curves correspond to minimal-spin contours with $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to spin-up lines with $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
 To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
 Refer to captionRefer to caption
-Figure 4: Contours of $\Delta M_{\rm min}$ and $B_{\rm max}$ for SSs (left) and NSs (right).
+Figure 5: Minimal-spin contours of $\Delta M$ and spin-up lines of $B$ for SSs (left) and NSs (right).
 The horizontal ordinate denotes the final mass after accretion.
-Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
+Solid curves correspond to $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
 Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
 The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.5.
-Although the curves of $\Delta M_{\rm min}$ for SSs are not significantly different from that for NSs, the differences between the contours of $B_{\rm max}$ for SSs and that for NSs are visible.
+Although the minimal-spin contours of $\Delta M$ for SSs are not significantly different from that for NSs, the differences between the contours of $B$, i.e.
+the spin-up lines, for SSs and that for NSs are visible.
 The dashed curves in Fig.
 5 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs.
-The constraints on EoS models are more significant for masses below about $1.6M_{\odot}$ than that for higher masses.
+The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$ than that for higher masses.
+If we can find some fast spinning pulsars with low masses, and if we can additionally know the lower limit of $B$, i.e.
+$B>B_{0}$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
 For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could.
-And on the other hand, if the magnetic field is higher than $8\times 10^{8}$ G, then an SS with mass approaching $2.2M_{\odot}$ could not be spun up to $P<4$ ms, but an NS could.
-As more and more masses of MSPs with increasing accuracy have been being measured, we can study the recycling of pulsars by comparison with observations in $MP$ diagram.
-Although from the data plotted in Fig.
-4 we could not yet disfavor neither SS nor NS models, finding more MSPs with $P<10$ ms with neither low mass or high mass in the future could help to put more strict constraints on the EoS of pulsars.
+As we mentioned before, each pulsar plotted in Fig.5 is above its spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using its data of $P$ and $\dot{P}$, so at this point if we take the lower limit magnetic field strength $B_{0}$ as $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ we could not yet disfavor neither SS nor NS models.
+Certainly, finding more MSPs of $P<10$ ms with low mass and more accurate estimation of $B_{0}$ in the future could help to put more strict constraints on the EoS of pulsars.
 5 Conclusions and discussions
@@
 Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs.
-The allowed regions in $MP$ diagram derived by the lower limit for $B$ and the upper limit for $\Delta M$ indicate that, it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
-Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled MSPs with neither low mass (below about $1.5M_{\odot}$) or very high mass (above about $2.2M_{\odot}$) in the future could be helpful.
+Under the same magnetic field strength, an SS could spin faster than an NS in the recycling process, and NSs would spin faster than SSs with the same amount of accreted mass.
+The minimal-spin contours show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same, but the differences between the spin-up lines for SSs and that for NSs are obvious.
+The magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs, so it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
+The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$, when the upper limit of the allowed magnetic field strength of an SS is higher than that of an NS.
+To give a more general and model-independent investigation, in this paper we keep the magnetic field strength $B$ as a parameter instead of replacing it in some model-dependent ways.
+It is worth noting that an accreting pulsar would have an lower limit of magnetic field strength (e.g.
+see Zhang & Kojima (2006)), so the constraints on NS model would be more strict than that on SS model.
+For the fast spinning pulsars with low masses, if we can additionally derive the lower limit of magnetic field strength theoretically from the magnetic field decay models, or if we can directly detect the magnetic field strength from the cyclotron absorption lines, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
 Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km.
 We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs.
-It is worth adding the role of EoS to other effects, such as the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion, to study the recycling of MSPs more comprehensively.
+To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as the pattern of disk-fed accretion flow and the magnetic field configuration outside the accreting star, which lead some simplifications in equations (1) and (2).
+In addition we do not consider the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion.
+Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
 The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions.
@@
 We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star.
-Finding more MSPs with spin periods less than 10 milliseconds and with accurate mass-measurement in the future, e.g.
-by SKA and FAST, could help to put more strict constraints on the nature of pulsars.
+Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on magnetic field strength in the future, e.g.
+by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.
 This work is supported by the National SKA Program of China (No. 2020SKA0120300) and the Young Top-notch Talent Cultivation Program of Hubei Province.

### v3

@@
 We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817.
-Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength.
-The magnetic field strength will be important in impeding the spin-up of NSs, especially to the ones with low masses, but the impeding effect would not be important to SSs.
-The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS.
-Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on magnetic field strength, could help to put more strict constraints on the EoS of pulsars.
-dense matter — stars: neutron — pulsars: general
+Combining the spin-up condition and the transferred angular momentum, and taking into account the evolution of magnetic field strength during accretion, we can constrain the spin-period and mass of an MSP.
+Our results show that the impeding effect of magnetic field on the spin-up of MSPs would be more significant for NSs than for SSs, especially for the ones with low masses.
+In the low-mass ($M$ below or around about $1.2M_{\odot}$) case, an SS can spin faster than an NS of the same mass by accreting the same amount of mass.
+Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on the amount of accreted mass, could help to put more strict constraints on the EoS of pulsars.
+keywords:
+dense matter – pulsars: general
+pubyear: 2015pagerange: The role of equation of state on the spin-up of millisecond pulsarsThe role of equation of state on the spin-up of millisecond pulsars
 1 Introduction
@@
 The strangeon star (SS) model was also proposed (Xu, 2003) to understand different manifestations of pulsars, where the strong coupling between quarks has grouped quarks into strangeons (i.e.
-“strange nucleons”) , and has passed various observational tests (see Lai et al.
-(2023) and references therein).
+“strange nucleons”) , and has passed various observational tests (see Lai et al., 2023, and references therein)
 The EoS of pulsars is essentially depends on the behavior of quantum chromodynamics (QCD) at low-energy scales, which is still a challenge for us to understand.
@@
 Both QSs and SSs are self-bound by strong interaction so that the surface density could still above the nuclear matter density, and their radii increase with increasing masses unless the mass approaches its maximum value.
-The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.5M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not.
+The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.4M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not.
 Therefore, testing whether pulsars are gravity-bound or self-bound provides a relatively less unambiguous probe to the nature of pulsars.
@@
 MSPs evolve from pulsars in binaries which have underwent the spin-up process, i.e.
-recycling precess, due to accreting mass from their companions.
-The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000, 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star.
+recycling process, due to accreting mass from their companions.
+The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000; Tauris et al., 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star.
 The observed differences of the spin distributions of recycled pulsars with different types of companions have been explained by the amount of accreted mass needed for a pulsar to achieve its equilibrium spin (Tauris et al., 2012).
 In addition, the magnetic field evolution of accreting pulsars would also affect the formation of MSPs.
-The spin-period evolutions of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006).
+The spin-period evolution of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006).
 The minimum spin-period of millisecond pulsars also depends the correlation between the accretion rate and the surface dipole magnetic field in LMXBs (Ertan & Alpar, 2021).
@@
 These studies discuss many parameters which do affect the location of the spin-up line in the $P\dot{P}$-diagram, e.g.
-the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979a).
+the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979b).
 It is worth noting that, the spin evolution of a pulsar also depends on the EoS which determines the pulsar’s global structure such as mass, radius and the moment of inertia.
 In this paper we will investigate the role of equation of state on the spin-up of MSPs.
-Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs.
+The EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817.
+In this paper we use two kinds of EoS models for strangeon stars (SSs): LX3630 (Lai & Xu, 2009; Gao et al., 2022) and Z-2023 (hybrid SSs) (Zhang et al., 2023).
+We also use two kinds of EoS models for neutron stars (NSs): AP4 (Akmal & Pandharipande, 1997) and X-2024 (hybrid NSs) (Xia, 2024).
+A hybrid SS is an SS which develops a strange quark matter core, so it is self-bound.
+A hybrid NS is an NS which develops a strange quark matter core, so it is gravity-bound.
+Therefore, we choose SSs in both LX3630 and Z-2023 models to represent self-bound SSs, and choose NSs in both AP4 and X-2024 to represent gravity-bound NSs.
+Considering the large error of NICER measurements (e.g.
+Miller et al.
+(2021)), we do not use their results to constrain EoS models.
+From the mass ($M$)-radius ($R$) curves shown in Fig.1, we can see that for the mass below $\sim 1.4M_{\odot}$ SSs would be more compact than NSs with the same mass.
+SSs could have very low masses and small radii, and the discrepancy between the results of LX3630 and AP4 are more obvious.
+We try to demonstrate that, for the spin-up process of MSPs there could be observational differences between SSs and NSs in the low-mass case.
+Figure 1: $M$-$R$ curves for SSs in LX3630 (red solid lines), hybrid SSs in Z-2023 (red dashed line), NSs in AP4 (blue solid line), and hybrid NSs in X-2024 (blue dashed line).
+In this paper we use the former two models to describe SSs and the latter two models to describe NSs
 From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process.
-To take into account the effect of accretion process, in Sec.3 we derive the minimal spin-period by assuming that all the angular momentum of the accreted material is added the the pulsar.
-Although the results of Sec.3 show that NSs would spin faster than SSs with the same amount of accreted mass and the same magnetic field, the minimal-spin contours derived in Sec.4 show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same.
-The results shown in Sec.4 indicate that the constraints on EoS models are significant for masses below about $1.5M_{\odot}$, especially when we get more accurate estimation of the magnetic field strength.
+To investigate the effect of accretion process, in Sec.3 we derive the spin-period as the function of accreted mass, taking into account the evolution of magnetic field strength during accretion.
+Then we can constrain the spin-period and mass of an MSP.
+The $M$-$P$ lines shown in Sec.4 indicate that the constraints on EoS models could be better for masses below or around about $1.2M_{\odot}$, especially when we could get more accurate constraints on the amount of accreted mass.
 Conclusions and discussions are made in Sec.5.
@@
 […]
-Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979a), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter.
-To exert a positive accretion torque, $\omega_{\rm c}$ is about 0.2-1 (Ghosh & Lamb, 1979b; Wang et al., 2011).
+Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979b), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter.
+To exert a positive accretion torque $\omega_{\rm c}$ should be about 0.2-1 (Ghosh & Lamb, 1979a), and the spin frequency of 716 Hz could indicate that $\omega_{\rm c}=0.55$ (Wang et al., 2011).
 The spin-up condition can therefore be expressed as the relation between $r_{\rm A}$ and $r_{\rm co}$, where the corotation radius $r_{\rm co}=[GMP^{2}/(4\pi^{2})]^{1/3}$ is defined as the radius where the Keplerian frequency matches the spin frequency of the pulsar.
-Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as $r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}$.
+Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as
+$r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}.$ (1)
 Because we will focus on the EoS-dependent results of the spin-up process, the dependence on other factors, such as truncation of the accretion disc and the magnetic field configuration outside the star, will not be included when we are comparing different EoS models.
-Taking into account the possible values for the factors $\phi$ and $\omega_{\rm c}$ mentioned before, we assume $\omega_{\rm c}^{2/3}/\phi=0.5$, so the spin-up condition is
-$r_{\rm A}\leq 0.5\ r_{\rm co}.$ (1)
+In the following we adopt $\phi=0.7$ and $\omega_{\rm c}=0.55$.
 The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar).
@@
 For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006).
-To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities.
-Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram.
+More importantly, the value of $B$ calculated from $P$ and $\dot{P}$ under any spin-down torque model should be very different from that in the accreting process which we will consider in Sec.3.
+Therefore, we keep $B$ as a parameter and plot the spin-up line in $MP$ diagram.
+Then assuming the mass accretion rate $\dot{M}=0.1\dot{M}_{\rm Edd}$ (the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 2\times 10^{-8}M_{\odot}\rm yr^{-1}$), and using equation (1), we can plot the spin-up line in $MP$ diagram.
 The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent.
-Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.1 shows the results for NS and SS models.
-Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively.
-We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass.
-The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
-Figure 1: Spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
-3 The minimal spin-period by accretion
-By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by ${\rm d}J=n\xi\sqrt{GMr_{\rm A}}{\rm d}M$, where $\xi\simeq 1$ is a numerical factor depending on the flow pattern (Ghosh & Lamb, 1979a) and $n$ is a dimensionless torque which could be assumed to be $\simeq 1$ (Tauris et al., 2012).
-Then the final spin-period $P$ of a recycled pulsar (from rest) from the initial mass $M_{1}$ to the final mass $M_{2}$ could be estimated by
-$P^{-1}_{\rm min}=\frac{1}{2\pi I}\int_{M_{1}}^{M_{2}}\sqrt{GMr_{\rm A}}{\rm d}M,$ (2)
-where $I$ is the (final) moment of inertia.
-We have assumed above that all the orbital angular momentum of the accreted material is transferred to the spin angular momentum of the pulsar, so $P$ in equation (2) is the minimal value of the final spin-period.
-In the following the values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms.
-Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables.
-To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumption as in Tauris et al.
-(2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium (but this assumption can be relaxed in deriving the minimal-spin contour which contains different values of $B$, see Sec.4).
-In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak.
-Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$.
-Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
-The $P_{\rm min}$ and $M$ (final mass) relation for SS and NS models are shown in Fig.2, denoted respectively by red lines and blue lines, where solid lines represent the case of $\Delta M=0.05M_{\odot}$ and dashed lines represent the case of $\Delta M=0.1M_{\odot}$.
-We can see that, with the same amount of accreted mass, NSs would spin faster than SSs.
-This is understandable: the radius of an NS is usually larger than that of an SS with the same mass, so an NS usually has larger $r_{\rm A}$ than an SSs, which means that an NS could gain more angular momentum than an SS when the accreted mass is the same.
-In Fig.2 we choose $B=5\times 10^{8}$ G, and smaller values of $B$ will lift the lines upwards, since smaller $r_{\rm A}$ will lead to smaller angular momentum transfer.
-Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G.
-The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively.
-In $MP$ diagram, the horizontal ordinate denotes the final mass after accretion.
+Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.2 shows the results for SSs (including LX3630 and Z-2023 models) and NSs (including AP4 and X-2024 models).
+The four spin-up lines of $B=5\times 10^{8}$ G correspond to SSs in LX3630 (red solid line), SSs in Z-2023 (red dashed line), NSs in AP4 (blue solid line), and NSs in X-2024 (blue dashed line), respectively.
+We can see that the spin-up lines of NSs (both AP4 and X-2024 models) are decreasing with the mass, and that of SSs (both LX3630 and Z-2023 models) are nearly independent of the mass.
+Figure 2: Spin-up lines of $B=5\times 10^{8}$ G for SSs in LX3630 (red solid line) and Z-2023 (red dashed line), and for NSs in AP4 (blue solid line) and X-2024 (blue dashed line).
+The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.2 indicates that by a sufficient mass-transfer an SS could have smaller values of $P$ than an NS in the low-mass case.
+The dashed lines in Fig.
+2 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with masses.
+When below about $M<1.4M_{\odot}$, the impeding effect would be less important for SSs than for NSs with the same values of $B$.
+This is understandable: with the same mass below $1.4M_{\odot}$, the radius of an NS is usually larger than that of an SS, i.e.
+an NS usually has larger $r_{\rm A}$ than an SS, which means that in the low-mass case it is more difficult to satisfy the spin-up condition (1) for an NS than that for an SS.
+The difference between SSs in LX3630 (red solid line) and NSs in AP4 (blue solid line) is more pronounced than that between SSs in Z-2023 (red dashed line) and NSs in X-2024 (blue dashed line).
+This can be inferred from Fig.1, which shows that at low masses, the discrepancy of $M$-$R$ curves between LX3630 and AP4 models is more obvious than that between Z-2023 and X-2024 models.
+3 Spin-periods after accretion
+By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by
+${\rm d}J=n\sqrt{GMr_{\rm m}}{\rm d}M,$ (2)
+where $n$ is a dimensionless torque.
+In principle, we can derive the final spin-period of a pulsar (from rest) after accreting the amount of mass $\Delta M=M-M_{0}$ (if the spin-up condition is satisfied).
+When considering the accretion process, two factors should be taken into account.
+(1) The torque $n$ depends on fastness parameter ($\Omega/\Omega_{\rm K}(r_{\rm m})$).
+We use the formula $n=1-\Omega/\Omega_{\rm K}(r_{\rm m})$ from Menou et al.
+(1999).
+This indicate that the torque decreases as the spin-up proceeds.
+(2) The evolution of magnetic field $B$ during accretion.
+The accretion leads to decay of magnetic fields, which then changes the magnetospheric radius $r_{\rm m}$.
+We use the relation between $B$ and the accreted mass $\delta M$ given by Zhang & Kojima (2006), $B\simeq 0.8B_{\rm f}(\delta M/M_{\rm cr})^{-1.75}$, where $B_{\rm f}=10^{8}$ G and $M_{\rm cr}=0.2M_{\odot}$.
+The spin-period with the amount of accreted mass $\Delta M$ derived from equation (2) is also EoS-dependent.
+The angular momentum $J=I\Omega$, where $I$ is the moment of inertia.
+The values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms.
+Then by assuming $\phi=0.7$ and $\dot{M}=0.1\dot{M}_{\rm Edd}$ as in Sec.2, equation (2) is related to $M$, $R$ and $I$.
+Given an EoS, the spin-period $P$ with the amount of accreted mass $\Delta M$ could be derived.
+The curves of spin-period $P$ after accretion of the amount of mass $\Delta M=0.1M_{\odot}$ are shown in Fig.3, where the horizontal ordinate denotes the final mass after accretion.
+The results of SSs in LX3630 and Z-2023 models are represented respectively by red solid line and red dashed line, and the results of NSs in AP4 and X-2024 models are represented respectively by blue solid line and blue dashed line.
+We can see that, with the same amount of accreted mass, in the low-mass case SSs would spin faster than NSs.
+Larger/smaller values of $\Delta M$ will move the lines downwards/upwards.
+Figure 3: The curves of $P$ after accretion of the amount of mass $\Delta M=0.1M_{\odot}$, as the function of the final mass $M$.
+The results of SSs in LX3630 and Z-2023 models are shown respectively by red solid line and red dashed line, and the results of NSs in AP4 and X-2024 models are shown respectively by blue solid line and blue dashed line.
+The fastness-dependent torque $n$ indicates that the spin-up itself will impede the further spin-up.
+The difference in tendencies of $P$ towards increasing mass between NSs and SSs shown in Fig.3 reflects the difference in such kind of impediment for NSs and SSs.
+When $M<1.4M_{\odot}$, an NS usually has a larger $r_{\rm m}$ than that of an SS with the same mass, which means that in the low-mass case the impediment of spin-up is more significant for NSs than that for SSs.
+This difference essentially reflects the distinction between stellar structures of gravity-bound NSs and self-bound SSs.
 4 Comparison with observations
 The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram.
-As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$.
+As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$ to derive the spin-up lines in $MP$ diagram, and take into account the evolution of $B$ in the spin-up process.
 Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
-If the values of both $B$ and $\Delta M$ of a recycled pulsar is known, we can plot the allowed region in $MP$ diagram which satisfies both equations (1) and (2), i.e.
-by putting Fig.1 and Fig.2 together.
-We show the allowed regions of SSs (left) and NSs (right) for $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G in Fig.3, where dashed line denotes the general spin-up line (cf.
-Fig.1) and solid line shows the minimal spin-period by accretion (cf.
-Fig.2).
-From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
-Refer to captionRefer to caption
-Figure 3: The allowed regions of SSs (left) and NSs (right) in $MP$ diagram, where solid curves denote minimal-spin lines with $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G, and dashed curves denote spin-up lines with $B=5\times 10^{8}$ G.
-However, in most cases we are not certain about both the value of $B$ and $\Delta M$ for recycled pulsars.
-If we estimate $B$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, then by setting $\sin\alpha=1$ we can get $B>1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$.
-It is interesting to noted that, each pulsar whose data of $M$ and $P$ is plotted in Fig.5 (see below) is above the spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue (https://www.atnf.csiro.au/research/pulsar/psrcat/(Manchester et al., 2005) under both SS and NS models.
-As mentioned before, in this paper we keep $B$ as a parameter, instead of replacing in some model-dependent ways.
-Then how to show the difference between SSs and NSs models by comparing with observations?
-Combining equations (1) and (2) we can get the minimal-spin contour by accretion of $\Delta M$, which is defined as the contour of $\Delta M$ in $MP$ diagram, representing the minimal spin-period by accretion of $\Delta M$ for different values of $B$.
-For a data point in $MP$ diagram, there exists both a minimal-spin contour and a spin-up line intersecting it.
-The value of $\Delta M$ of the minimal-spin contour is the minimal accreted mass of the pulsar, while the value of $B$ of the spin-up line is the upper limit on its magnetic field strength.
-If we can use some other methods to constrain $\Delta M$ (e.g.
-from the binary evolution) and $B$ (e.g.
-theoretically from the magnetic field decay in accretion, or observationally from the cyclotron absorption lines), it would be hopeful to see that whether this pulsar is more likely to be an NS or an SS.
-From Fig.3 we can infer that low mass and fast spinning pulsars would present significant different parameters between SS and NS models, so we use a hypothetical data point in $MP$ diagram as an example in Fig.4.
-If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G.
-One the other hand, if this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
-This example shows that the difference in the the upper limit on $B$ between SS and NS models is significant for low mass and fast spinning pulsars.
-Figure 4: We use a hypothetical data point in $MP$ diagram as an example.
-If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G.
-If this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
-To compare with actual data, we extend the curves in Fig.4 to different values of $B$ and $\Delta M$, and we show in Fig.5 the results for SSs (left) and NSs (right).
-Solid curves correspond to minimal-spin contours with $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to spin-up lines with $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
-To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
-Refer to captionRefer to caption
-Figure 5: Minimal-spin contours of $\Delta M$ and spin-up lines of $B$ for SSs (left) and NSs (right).
-The horizontal ordinate denotes the final mass after accretion.
-Solid curves correspond to $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B/10^{8}{\rm G}=2,4,8,12$, increasing upwards.
+The spin-up lines derived in Section 2 do not consider the supplied mass, and the spin-periods after accretion of $\Delta M$ derived in Section 3 do not consider the spin-up condition, so we can derive the spin-periods after accretion of $\Delta M$ by putting Fig.2 and Fig.3 together.
+In addition, because we consider the evolution of $B$ during accretion, the spin-up lines depending on the value of $B$ can be determined by the accreted mass.
+In the upper panel of Fig.4, the orange solid line shows the spin-periods of SSs in LX3630 model with mass $M$ after accretion of $\Delta M=0.1M_{\odot}$, and the orange dashed line is the spin-up line of SSs in LX3630 whose magnetic field strength equals to the final $B$ after accretion of $\Delta M=0.1M_{\odot}$.
+Therefore, when $\Delta M=0.1M_{\odot}$, the spin-period of an SS in LX3630 lies in the curve formed by taking the higher parts of these two lines.
+The red solid and dashed lines are the corresponding results of SSs in Z-2023 model.
+From the observational point of view, it would be easier to constrain the upper limit of $\Delta M$, then we can get the allowed region which lies above both solid and dashed lines.
+We can see that the allowed region for SSs (covered by both orange and red patches) in LX3630 is bigger than and encloses that for Z-2023 (covered by red patch).
+The lower panel of Fig.4 shows the same lines as in the upper one, for the results of NSs in AP4 model (blue solid and dashed lines) and X-2024 model (green solid and dashed lines).
+The allowed region for X-2024 model (covered by both green and blue patches) is bigger than and encloses that for AP4 model (covered by blue patch).
+Comparing the upper and lower panels of Fig.4 we can see that, the allowed region of SSs (in both LX3630 and Z-2023 models) covers the lower left region in $MP$ diagram where NSs (in both AP4 and X-2024 models) cannot reach, indicating that it could be easier for an SS to be spun-up to a period less than 3 ms than that for an NS, when the mass is below about $1.2M_{\odot}$.
+Figure 4: The allowed regions of SSs (upper) and NSs (lower) in $MP$ diagram.
+The upper panel shows the spin-periods of SSs with mass $M$ after accretion of $\Delta M=0.1M_{\odot}$ in LX3630 (orange solid line) and Z-2023 (red solid line), and the spin-up line of SSs whose magnetic field strength equals to the final $B$ after accretion of $\Delta M=0.1M_{\odot}$ in LX3630 (orange dashed line) and Z-2023 (red dashed line).
+If $\Delta M<0.1M_{\odot}$, the allowed region for SSs in LX3630 is covered by both orange and red patches, and that for SSs in Z-2023 is covered by red patch.
+The lower panel of shows the same results as in the upper one, for NSs in AP4 model (blue solid line, blue dashed line, and blue patch) and in X-2024 model (green solid line, green dashed line, and patches in both blue and green).
+As $\Delta M$ increases/decreases, all the lines for $P$ after accretion of $\Delta M$ (solid lines) and spin-up lines (dashed lines) in Fig.4 will shift downwards/upwards.
+The allowed region of each model will be above both the solid and dashed lines in the corresponding color.
+For some values of $\Delta M$, solid line will be totally above or below the dashed line.
+In all cases, for a certain value of $\Delta M$ the spin-period of an MSP lies in the curve formed by taking the higher parts of solid and dashed lines in the same color, and for an upper limit of $\Delta M$ the allowed region of an MSP lies above both solid and dashed lines in the same color.
+The above example shows that, the role of EoS on the spin-up of MSPs could be manifested in low-mass (below about $1.2M_{\odot}$) case.
+As also shown in Fig.2 and Fig.3, the difference between SSs in LX3630 and NSs in AP4 is more pronounced than that between SSs in Z-2023 and NSs in X-2024 model.
+To compare with observational data, we extend the curves in Fig.4 to different values of $\Delta M$, and we show in Fig.5 the results for SSs (upper) and NSs (lower).
+Each line is derived by taking the higher parts of line of $P$ after accretion of $\Delta M$ and line of spin-up, for a certain value of $\Delta M$.
+In the upper panel, the results for SSs in SS model (LX3630) and hybrid SS model (Z-2023) are shown by red solid line and red dashed line, respectively.
+In the lower panel, the results for NSs in NS model (AP4) and hybrid NS model (X-2024) are shown by blue solid line and blue dashed line, respectively.
+In both panels, $\Delta M/M_{\odot}=0.04,0.06,0.1,0.2$, increasing downwards.
+We choose the masses of pulsars whose $P<15$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
+Figure 5: $M$-$P$ curves after accretion of $\Delta M$, for SSs in LX3630 (red solid line in upper panel), SSs in Z-2023 (red dashed line in the upper panel), NSs in AP4 (blue solid line in lower panel), and NSs in X-2024 (blue dashed line in lower panel).
+For each model, we show the results for $\Delta M/M_{\odot}=0.04,0.06,0.1,0.2$, increasing downwards.
 Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
-The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.5.
-Although the minimal-spin contours of $\Delta M$ for SSs are not significantly different from that for NSs, the differences between the contours of $B$, i.e.
-the spin-up lines, for SSs and that for NSs are visible.
-The dashed curves in Fig.
-5 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs.
-The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$ than that for higher masses.
-If we can find some fast spinning pulsars with low masses, and if we can additionally know the lower limit of $B$, i.e.
-$B>B_{0}$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
-For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could.
-As we mentioned before, each pulsar plotted in Fig.5 is above its spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using its data of $P$ and $\dot{P}$, so at this point if we take the lower limit magnetic field strength $B_{0}$ as $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ we could not yet disfavor neither SS nor NS models.
-Certainly, finding more MSPs of $P<10$ ms with low mass and more accurate estimation of $B_{0}$ in the future could help to put more strict constraints on the EoS of pulsars.
+The different effects on parameters of recycled MSPs between SSs and NSs can be seen from Fig.5.
+The discrepancy between SSs in LX3630 (red solid line) and NSs in AP4 (blue solid line) is bigger than that between SSs in Z-2023 (red dashed line) and NSs in X-2024 (blue dashed line).
+In the low-mass ($M$ below about $1.2M_{\odot}$) case, an SS can spin faster than an NS of the same mass by accreting the same amount of mass $\Delta M$, and the difference is more significant when $\Delta M<0.1M_{\odot}$.
+The constraints on EoS models could be better for masses below about $1.2M_{\odot}$ than that for higher masses.
+If we can find some fast spinning pulsars with low masses and additionally know the upper limit of $\Delta M$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
+For example, if we can constrain that $\Delta M<0.04M_{\odot}$, then an NS with $M<1.2M_{\odot}$ could not be spun up to $P<11$ ms, but an SS could.
+Furthermore, we can see that if we only consider SSs in LX3630 model, the discrepancy between SSs and NSs would be more significant.
+Certainly, finding more MSPs of $P<15$ ms with low mass and more accurate estimation of $\Delta M$ in the future could help to put more strict constraints on the EoS of pulsars.
 5 Conclusions and discussions
 The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries.
-We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model.
-Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs.
-Under the same magnetic field strength, an SS could spin faster than an NS in the recycling process, and NSs would spin faster than SSs with the same amount of accreted mass.
-The minimal-spin contours show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same, but the differences between the spin-up lines for SSs and that for NSs are obvious.
-The magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs, so it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
-The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$, when the upper limit of the allowed magnetic field strength of an SS is higher than that of an NS.
-To give a more general and model-independent investigation, in this paper we keep the magnetic field strength $B$ as a parameter instead of replacing it in some model-dependent ways.
-It is worth noting that an accreting pulsar would have an lower limit of magnetic field strength (e.g.
-see Zhang & Kojima (2006)), so the constraints on NS model would be more strict than that on SS model.
-For the fast spinning pulsars with low masses, if we can additionally derive the lower limit of magnetic field strength theoretically from the magnetic field decay models, or if we can directly detect the magnetic field strength from the cyclotron absorption lines, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
+We investigate the spin-up of MSPs of NSs (in AP4 and X-2024 models) and SSs (in LX3630 and Z-2023 models).
+Combining the spin-up condition and the transferred angular momentum, we can constrain the spin-period and mass of an MSP.
+When considering the accretion process, we take into account the evolution of magnetic field during accretion and the fastness-dependent torque.
+Our results show that, for the spin-up process of MSPs there could be observable differences between SSs and NSs.
+The impeding effect of magnetic field on the spin-up of MSPs would be more significant for NSs than for SSs, especially for the ones with masses below about $1.4M_{\odot}$.
+By accreting the same amount of mass, an SS with mass below about $1.2M_{\odot}$ can spin faster than an NS of the same mass, implying that it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
+Therefore, our results show that the constraints on EoS models would be better in such low-mass case.
+For the fast spinning pulsars with low masses, if we can additionally constrain the amount of accreted mass $\Delta M$ (which could be inferred by the nature of donor stars), it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
 Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km.
 We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs.
-To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as the pattern of disk-fed accretion flow and the magnetic field configuration outside the accreting star, which lead some simplifications in equations (1) and (2).
-In addition we do not consider the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion.
+To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as $\phi$ determined by the pattern of disk-fed accretion flow and the critical fastness $\omega_{\rm c}$ determining the accretion torque.
+Different choices of $\phi$ and $\omega_{\rm c}$ will influence the quantitative results, shifting the curves in Figs.2, 3, 4, and 5 upwards or downwards, but the qualitative conclusion that an SS could be more easily to form a low-mass and fully recycled MSP than an NS will not change.
+In addition, the expressions of the fastness-dependent torque and the evolution of magnetic field during accretion are also model-dependent.
 Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
-The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions.
-A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value.
-This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars.
+The observable differences are manifested in the low-mass case, which can be inferred from the fact that the discrepancy of $M$-$R$ curves between the two SS models and the two NS models is more significant in such low-mass case.
+Following the same reasoning, the difference between SSs in LX3630 and NSs in AP4 is more pronounced than that between SSs in Z-2023 and NSs in X-2024.
+Although the differences between the two SS models and the two NS models are obvious enough, if we take LX3630 model and AP4 model to represent EoS models of gravity-bound and self-bound respectively, we could get more robust conclusions.
+A gravity-bound pulsar, e.g.
+shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, e.g.
+shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value.
+This difference leads to different evolutions of $r_{\rm m}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars.
 We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star.
-Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on magnetic field strength in the future, e.g.
+Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on the amount of accreted mass in the future, e.g.
 by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.
-This work is supported by the National SKA Program of China (No. 2020SKA0120300) and the Young Top-notch Talent Cultivation Program of Hubei Province.
+Acknowledgements
+We would like to thank Prof.
+Renxin Xu in PKU, Prof.
+Wencong Chen in QUT, Prof.
+Chengjun Xia in YZU and Dr.
+Chen Zhang in HKUST for useful discussions, and to thank an anonymous referee for useful comments and suggestions.
+This work is supported by the National SKA Program of China (No.
+2020SKA0120300).
+Data Availability
+The data underlying the work in this paper are available upon reasonable request.